OpenStudy (anonymous):

If the polygons are similar, which ratio shows the relationship of the area of the larger polygon to the area of the smaller polygon? the bigger polygon has a measurement of 12 and the smaller polygon has a measuremtent of 6 and 8. 4/1 9/4 3/2

OpenStudy (anonymous):

is there a picture to go with this?

OpenStudy (anonymous):

http://i43.tinypic.com/34spb9u.jpg

OpenStudy (anonymous):

@completidiot here is the picture

OpenStudy (anonymous):

you need to use proportions in order to determine the missing side area of a trapezoid - A= 1/2 (b1+b2)(h) ratio A1/A2 = (b11+b21)/(b21+b22)

OpenStudy (anonymous):

On the ratio where did you get 11 and 21

OpenStudy (anonymous):

those arent numbers those are part of the variable

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i was trying to differentiate the variables by using numbers

OpenStudy (anonymous):

but then how do you find the missing side

OpenStudy (anonymous):

b1 - base on top of trapezoid b2- base on the bottom b11- base on top of trapezoid 1 b21- base on bottom of trapezoid 1

OpenStudy (anonymous):

we know that the 2 trapezoids are similar so we can use a proportion to determine the missing side

OpenStudy (anonymous):

8 is proportional to the missing side and 6 is proportional to 12

OpenStudy (anonymous):

okay so then it would be 12/6=2 so then 8x2 = 16 so then that is the missing side

OpenStudy (anonymous):

it can be set up as 8/6= x/12 where x is the missing side

OpenStudy (anonymous):

yes pretty much

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

But then what would be height

OpenStudy (anonymous):

well, when you're doing the ratio, you will realize that the height cancels out so it wont be needed

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

But wait then that will give you 2/1 that is not a choice

OpenStudy (anonymous):

yeah I know the area of the smaller one is 7 then the bigger one is 14.... 14/7 it gives you 2/1 when simplified

OpenStudy (anonymous):

oh wait sorry, i forgot, height is also proportional

OpenStudy (anonymous):

height of 2nd one would be 2h height of 1st one would be h

OpenStudy (anonymous):

SORRY

OpenStudy (anonymous):

its okay LOL

OpenStudy (anonymous):

But then how do you find h on the first one

OpenStudy (anonymous):

again, the h cancels out

OpenStudy (anonymous):

okay so then 2 is for the 12 or the 6 and since I have you here can you help me out on this please A bakery charges $0.10 for a cookie that is 2 inches in diameter. If the price is proportional to the area, how much do they charge for an extra large cookie that is 10 inches in diameter? Assume that each cookie is shaped of a circle.

OpenStudy (anonymous):

Thanks I got 4/1